OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = 1/3x − 3 y = 1/3x + 17 y = −3x − 3 y = −3x + 17

OpenStudy (johnweldon1993):

we have a point....(6,-1) we have an equation....and we want a perpendicular line right? a perpendicular line has the negative reciprocal slope of the other line ...so 3x + y = 7 in slope intercept form is y = -3x + 7 the negative reciprocal of -3 is? well -3...is the same as -3/1 right? the negative (turns this into positive) reciprocal (flip the fraction) is 1/3x so we have a slope now....and a point....we need the point slope formula y - y1 = m(x-x1) m is slope (6, −1) is our point (x1 , y1) plug everything in..what do you get?

OpenStudy (anonymous):

c?

OpenStudy (johnweldon1993):

y - (-1) = 1/3(x-6) y + 1 = 1/3(x-6) = 1/3x -2 y = 1/3x -3